A plane is flying within sight of the Gateway Arch in St. Louis, Missouri, at an elevation of 40,000 ft. The pilot would like to estimate her distance from the Gateway Arch. She finds that the angle of depression to a point on the ground below the arch is 19°. (Round your answers to the nearest foot.)
(a) What is the distance between the plane and the base of the arch?
(b) What is the distance between a point on the ground directly below the plane and the base of the arch?

Answer by mananth(16946)


Sin 19 = 40000/distance
distance from plane to the arch = 40000/sin(19
=122862.139 ft
tan 19 = 40000/x
x= 40000/tan 19
= 116168.435